# Generalizable and Relightable Gaussian Splatting for Human Novel View Synthesis

### v1

@@ Appendix A Preliminary
3DGS [21] is an explicit point-based representation that models a scene as a set of 3D Gaussians. Each Gaussian is parameterized by its position $\mathbf{p}$, a covariance matrix $\mathbf{\Sigma}$ (constructed from a rotation vector $\mathbf{r}$ and a scale vector $\mathbf{s}$), opacity $\alpha$, and color $\mathbf{c}$. In our framework, the color $\mathbf{c}$ can be generalized to represent any attribute that needs to be projected into image space for 2D-to-3D supervision. Then each gaussian can be expressed as:
$f(\mathbf{x}|\mathbf{p},\mathbf{\Sigma})=\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{p})^{\top}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{p})\right),$ (15)
@@
$\mathbf{C}=\sum_{i=1}^{N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})\mathbf{c}_{i},$ (16)
where $\mathbf{c}_{i}$ is the color of each 2D Gaussian, and $\alpha_{i}$ is the blending weight derived from the learned opacity and 2D Gaussian distribution [65].
@@
The whole BRDF [5] employed in our method is composed of a diffuse term $f_{\text{d}}$ and a specular term $f_{\text{s}}$, defined as:
$f\left(\bm{\omega}_{i},\bm{\omega}_{o},\mathbf{a},\gamma\right)=\underbrace{\frac{\mathbf{a}}{\pi}}_{f_{d}}+\underbrace{\frac{D\left(\mathbf{h};\gamma\right)\cdot F\left(\bm{\omega}_{o},\mathbf{h}\right)\cdot G\left(\bm{\omega}_{i},\bm{\omega}_{o},\mathbf{h};\gamma\right)}{\left(\bm{\omega}_{i}\cdot\bm{n}\right)\cdot\left(\bm{\omega}_{o}\cdot\bm{n}\right)}}_{f_{s}}$ (17)
where $\mathbf{h}=\frac{\left(\bm{\omega}_{i}+\bm{\omega}_{o}\right)}{2}$ denotes the half-vector between the incoming direction $\bm{\omega}_{i}$ and outgoing direction $\bm{\omega}_{o}$. The functions $D\left(\cdot\right)$, $F\left(\cdot\right)$, and $G\left(\cdot\right)$ correspond to the normal distribution function (NDF), Fresnel term, and geometric term, respectively. The parameter $\mathbf{a}$ represents the albedo, while $\gamma$ denotes the roughness.
Appendix B Experiment Details
@@
We employ Cycles renderer in Blender to render the datasets, positioning 16 cameras uniformly in a circular arrangement with an angular interval of 22.5°between adjacent cameras. Each human scan is rendered from every input camera pair, along with three novel views sampled from the intersection arc between the input cameras, under five randomly selected HDR environment maps. For each rendering, we generate corresponding outputs, including albedo, normal, depth, foreground mask, ambient occlusion, indirect light, shading, and relighting images. Since the scans are captured under uniformly illuminated conditions, their texture maps are approximately treated as intrinsic albedo in this dataset. Our training set consists of 600 scans from the Twindom [46] dataset and 277 scans from the THuman2.0 [57] dataset. For evaluation, we reserve 200 scans from Twindom and 60 scans from THuman2.0 as the test set.
Ablation study.
Table 2 compares novel view synthesis (NVS) and depth accuracy between GPS-Gaussian [63] and our method, highlighting the effectiveness of our lighting-invariant priors in enhancing both geometric reconstruction and appearance fidelity. Note that we evaluate depth accuracy using End-Point Error (EPE) and the 1-pixel accuracy metric, which measure the average disparity error and the percentage of pixels with depth error less than one pixel, respectively.
Table B demonstrates the effectiveness of the proposed normal refinement module by evaluating surface normal quality using the MAE metric. While the incorporation of lighting-invariant priors facilitates the recovery of smooth normals, fine-grained geometric details remain lacking due to the limitations of convex upsampling. In contrast, our normal refinement module successfully captures high-frequency details, leading to improvements in MAE.
Table 2: NVS and Depth Evaluation
Model NVS Depth
@@
GPS-Gaussian [63] 30.444 0.913 0.107 1.463 68.91
Ours 31.694 0.949 0.067 0.692 85.39
Table 3: MAE Comparison
Method MAE $\downarrow$
GPS-Gaussian [63] 8.145
Coarse Normal $\mathbf{N}_{c}$ 5.903
Refined Normal $\mathbf{N}_{f}$ 5.369
@@
We conducted a user study to evaluate the visual plausibility of our method on the proposed test set. In a pairwise comparison setup between our method and each baseline, 50 participants (including 40 students specializing in computer vision and graphics and 10 individuals from the general public) were asked to select the more visually convincing relighting result from each image pair. The study was conducted using 50 image samples, and we report the preference rate, defined as the fraction of times participants favored our results over those of the baseline methods in Table 4.
@@
We present additional relighting results of various performers under diverse lighting conditions in Fig. 8. The photorealistic quality, lighting consistency, and physical plausibility observed in these results highlight the effectiveness and generalization capability of our proposed method.
@@
Figure 9: Dynamic relighting results under diverse lighting conditions and novel viewpoints.
Although GRGS is not explicitly designed with a temporal consistency module for dynamic scenarios, it can be extended to dynamic settings under uniform lighting conditions, as shown in Fig. 9. By treating input images as diffuse albedo, the method mitigates most flickering artifacts caused by temporal inconsistencies.
Appendix C Implementation Details
In the LGR module, we set the number of radiance feature scales $K=3$, with feature dimensions of 32, 48, and 96, respectively, for$\{\mathbf{F}_{s}\}_{s=1}^{K}$. In the PGNR module, the target environment map is convolved from a resolution of $1024\times 512$ down to $64\times 32$ using a shininess exponent of $l=16$. Owing to the efficient light parameterization, we sample only 40 rays per Gaussian point oriented around the surface normal $\bm{n}$ for PBR, enabling fast and high-quality shading under arbitrary environment maps. We train our GRGS framework using the AdamW [32] optimizer with an initial learning rate of $2e^{-4}$. For the LGR module, we use a batch size of 2, while a batch size of 1 is adopted for the PGNR module. Note that during PGNR training, the LGR module is jointly optimized to facilitate more accurate localization of Gaussian point positions.
Appendix D Optimization
@@
The depth loss $\mathcal{L}_{\text{depth}}$ is defined as the weighted L1 distance between the predicted depth sequence $\{\mathbf{d}_{1},\cdots,\mathbf{d}_{N}\}$ and ground truth depth $\mathbf{d}_{\text{GT}}$ with exponentially increasing weights following [29]:
$\mathcal{L}_{\text{depth}}=\sum_{i=1}^{N}\mu^{N-i}||\mathbf{d}_{i}-\mathbf{d}_{\text{GT}}||_{1}$ (19)
where $\mu$ is set to 0.9. For the normal loss, given the ground truth $\mathbf{N}_{\text{GT}}$, we combine the L1 distance and a perceptual loss [61] to ensure both geometric accuracy and surface smoothness:
@@
We define the overall loss as a combination of $\mathcal{L}_{\text{albedo}}$, material smoothness loss $\mathcal{L}_{\text{smooth}}$, light transport loss $\mathcal{L}_{\text{LT}}$, and PBR loss $\mathcal{L}_{\text{PBR}}$ to facilitate high-quality appearance reconstruction and relighting via 2D-to-3D supervision:
@@
Note that after PBR and rasterization, the material maps: $\mathbf{A},\bm{\Gamma}$; the visibility $V$ parameterized by SH map $\mathbf{Y}_{\text{vis}}$; the indirec light $\mathbf{L}_{\text{ind}}$ parameterized by SH map $\mathbf{Y}_{\text{ind}}$; the direct light $\mathbf{L}_{\text{d}}$, and the PBR color $\mathbf{C}_{\text{PBR}}$ are all projected into the 2D image space as $\mathbf{I}_{\text{albedo}}$, $\mathbf{I}_{\text{rough}}$, $\mathbf{I}_{\text{ao}}$, $\mathbf{I}_{\text{indl}}$, $\mathbf{I}_{\text{dl}}$, $\mathbf{I}_{\text{PBR}}$, respectively. Thus, they can be supervised via corresponding 2D ground truths. Specifically, the albedo loss is composed of an L1 loss and a perceptual loss to measure the difference between the predicted albedo and the ground truth $\mathbf{I}_{\text{albedo}}^{\prime}$:
@@
To promote spatially smooth material estimation, we adopt a bilateral smoothness loss following [12]:
$\mathcal{L}_{\text{smooth}}=\lambda_{3}||\nabla\mathbf{\mathbf{I}_{\text{albedo}}}||\text{exp}\left(-||\nabla\mathbf{I}_{\text{albedo}}^{\prime}||\right)+\lambda_{4}||\nabla\mathbf{\mathbf{I}_{\text{rough}}}||\text{exp}\left(-||\nabla\mathbf{I}_{\text{albedo}}^{\prime}||\right)$ (23)
Note that ground truth roughness maps are not available; thus, the roughness map $\bm{\Gamma}$ is learned implicitly through the entire framework. The light transport loss is combined with ambient occlusion loss $\mathcal{L}_{\text{ao}}$, direct light loss $\mathcal{L}_{\text{d}}$, and indirect light loss $\mathcal{L}_{\text{ind}}$, we use L1 distance to measure the predicted and ground truth ones:
@@
Appendix E Limitations
While our method delivers high-quality relighting results on human subjects, several challenging cases remain. Due to the inherent limitations of the adopted rendering equation, our approach is unable to accurately model transparent materials such as eyeglasses and struggles to handle extremely thin structures, including hair strands. Future endeavors could explore the implementation of advanced reflectance models to further enhance the realism of relighting.

### v2

@@ 6 Preliminary
3DGS [kerbl20233d] is an explicit point-based representation that models a scene as a set of 3D Gaussians. Each Gaussian is parameterized by its position $\mathbf{p}$, a covariance matrix $\mathbf{\Sigma}$ (constructed from a rotation vector $\mathbf{r}$ and a scale vector $\mathbf{s}$), opacity $\alpha$, and color $\mathbf{c}$. In our framework, the color $\mathbf{c}$ can be formulated to represent any attribute requiring optimization under 2D-to-3D supervision. Then each gaussian can be expressed as:
$f(\mathbf{x}|\mathbf{p},\mathbf{\Sigma})=\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{p})^{\top}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{p})\right)$ (15)
@@
$\mathbf{C}=\sum_{i=1}^{N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})\mathbf{c}_{i}$ (16)
where $\mathbf{c}_{i}$ is the color of each 2D Gaussian, and $\alpha_{i}$ is the blending weight derived from the learned opacity and 2D Gaussian distribution [zwicker2001surface].
@@
The whole BRDF [burley2012physically] employed in our method is composed of a diffuse term $f_{\text{d}}$ and a specular term $f_{\text{s}}$, defined as:
$f(\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o},\mathbf{a},\gamma)=\underbrace{\tfrac{\mathbf{a}}{\pi}}_{f_{d}}+\underbrace{\tfrac{D(\mathbf{h};\gamma)F(\boldsymbol{\omega}_{o},\mathbf{h})G(\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o},\mathbf{h};\gamma)}{(\boldsymbol{\omega}_{i}\!\cdot\!\boldsymbol{n})(\boldsymbol{\omega}_{o}\!\cdot\!\boldsymbol{n})}}_{f_{s}}$ (17)
where $\mathbf{h}=\frac{\left(\boldsymbol{\omega}_{i}+\boldsymbol{\omega}_{o}\right)}{2}$ denotes the half-vector between the incoming direction $\boldsymbol{\omega}_{i}$ and outgoing direction $\boldsymbol{\omega}_{o}$. The functions $D\left(\cdot\right)$, $F\left(\cdot\right)$, and $G\left(\cdot\right)$ correspond to the normal distribution function (NDF), Fresnel term, and geometric term, respectively. The parameter $\mathbf{a}$ represents the albedo, while $\gamma$ denotes the roughness.
Figure 8: Network architecture. For simplicity, we consider a single-view image $\hat{I}$ as the input, while $\bar{F}_{K}$ denotes the final-scale image feature extracted from the other view.
7 Implementation Details
@@
We construct our dataset following a similar multi-view setup as GPS-Gaussian [zheng2024gps], but unlike GPS-Gaussian, which only renders novel view RGB images, our data construction explicitly accounts for geometry, material, and illumination information. We employ Cycles renderer in Blender to render the datasets, positioning 16 cameras uniformly in a circular arrangement with an angular interval of 22.5° between adjacent cameras. Each human scan is rendered from every input camera pair, along with three novel views sampled from the intersection arc between the input cameras, under five randomly selected HDR environment maps. For each rendering, we generate corresponding outputs, including albedo, normal, depth, foreground mask, ambient occlusion, indirect light, shading, and relighting images. Since the scans are captured under uniformly illuminated conditions, their texture maps are approximately treated as intrinsic albedo in this dataset. Our training set consists of 600 scans from the Twindom [twindom] dataset and 277 scans from the THuman2.0 [yu2021function4d] dataset. For evaluation, we reserve 200 scans from Twindom and 60 scans from THuman2.0 as the test set.
Network Architecture and Hyperparameter.
@@
Table 3: Runtime & relighting comparisons between spherical harmonics-based visibility (SH-Vis) modeling and ray-traced visibility (RT-Vis) under various ray sampling configurations.
SH-Vis RT-Vis
Sampling No. 40 128 256 40 128 256
Runtime $\downarrow$ 49 71 103 230 819 1536
Relit PSNR $\uparrow$ 27.977 27.968 28.041 26.762 27.593 27.812
8 More Comparison Results
Computation efficiency.
To quantitatively validate the computational efficiency of our method, we conduct runtime and relighting comparisons between our spherical harmonics-based visibility (SH-Vis) modeling and ray-traced visibility (RT-Vis) proposed in R3DGS under various ray sampling configurations (40, 128, 256) as shown in Table 3. The results demonstrate that our SH-Vis design achieves a significant speedup (at least 4× faster), compared to ray-traced visibility, while also delivering better relighting quality. Although SH-Vis-256 yields slightly higher image quality, its runtime doubles compared to SH-Vis-40, which hinders real-time inference. Moreover, SH-Vis-40 achieves comparable PSNR due to effective light parametrization, making it a more practical choice for our experiments. In contrast, RT-Vis incurs substantially longer runtimes and yields lower relighting quality, primarily because Gaussian point clouds lack explicit surfaces, forcing ray tracing to rely on computationally expensive volumetric intersection, which is not yet as accurate or efficient as mesh-based methods. In comparison, SH-Vis can learn high-frequency details from ground truth supervision, enabling it to achieve more accurate and visually appealing relighting results.
Table 4: Runtime & relighting comparisons against 3D baselines.
Ours R3DGS ARGS
Sampling NO. 40 40 256 40 256
Runtime $\downarrow$ 49 221 1480 1103 6318
Relit PSNR $\uparrow$ 27.977 19.214 21.983 21.336 23.879
Furthermore, we conduct runtime and relighting quality comparisons against 3D baselines in Table 4, including R3DGS and ARGS. R3DGS accelerates rendering by precomputing visibility via ray tracing. To ensure a fair evaluation, we include this precomputation time in the total runtime. Specifically, we compare R3DGS and ARGS under two ray sampling configurations: 40 rays and 256 rays. The 256-ray setting is consistent with our main paper experiments, while the 40-ray setting is included to highlight efficiency under reduced sampling. Both R3DGS and ARGS compute visibility using ray tracing over 3D Gaussian point clouds. ARGS suffers from significantly slower rendering speed due to its heavy Style-Unet architecture. However, our method achieves superior relighting quality while being significantly more computationally efficient due to the SH-Vis and light parametrization.
@@ 8 More Comparison Results
We conducted a user study to evaluate the visual plausibility of our method on the proposed test set. In a pairwise comparison setup between our method and each baseline, 50 participants (including 40 students specializing in computer vision and graphics and 10 individuals from the general public) were asked to select the more visually convincing relighting result from each image pair. The study was conducted on 50 image samples. We report the preference rate (defined as the fraction of times participants favored our results over those of the baselines) in Table 5, where participants consistently rated our relighting outputs as more visually convincing and physically coherent than those of baseline methods.
@@
9 Further Analysis
Geometry reconstruction.
Table 6 demonstrates the effectiveness of the proposed normal refinement module by evaluating surface normal quality using the MAE metric. While the incorporation of lighting-robust features facilitates the recovery of smooth normals, fine-grained geometric details remain lacking due to the limitations of convex upsampling. In contrast, our normal refinement module successfully captures high-frequency details, leading to improvements in MAE.
Table 6: MAE Comparison
Method GPS-Gaussian [zheng2024gps] Coarse Normal Refined Normal
MAE $\downarrow$ 8.145 5.903 5.369
Inference speed
The end-to-end pipeline, including geometry reconstruction, material prediction, and relighting, runs at approximately 20 FPS on an RTX 4090 GPU. The runtime breakdown of each module is shown in Table 7.
@@ 9 Further Analysis
We present additional relighting results of various performers under diverse lighting conditions in Fig. 9. The photorealistic quality, lighting consistency, and physical plausibility observed in these results highlight the effectiveness and generalization of our proposed method.
Table 7: Run time breakdown of each module.
Methods Geometry Gaussian attributes Rendering Total
Run time (ms) 11 21 17 49
@@
Although GRGS is not explicitly designed with a temporal consistency module for dynamic scenarios, it can be extended to dynamic settings under uniform lighting conditions, as shown in  Fig. 10Fig. 11 and  Fig. 12. By treating input images as diffuse albedo, the method mitigates most flickering artifacts caused by temporal inconsistencies. We further provide a qualitative comparison on dynamic sequences against R3DGS [gao2024relightable] and ARGS [li2024animatable], as shown in Fig. 13 and Fig. 14. R3DGS struggles to reconstruct geometry from sparse-view inputs, resulting in severe artifacts in relighting. ARGS alleviates this issue by employing a body template, which improves geometric consistency. However, its template deformation capability is limited, making it difficult to handle loose or highly deformable clothing such as long dresses. Moreover, both approaches require extensive optimization, taking 3–4 days to process a 200-frame sequence, which restricts practicality.
In contrast, our method requires no extra optimization, achieves near real-time inference, and does not rely on body templates, enabling robust generalization to diverse human identities and clothing styles. With the proposed PGNR module, our approach further delivers more natural, coherent, and physically plausible relighting results across dynamic sequences.
10 Loss Functions
@@
The depth loss $\mathcal{L}_{\text{depth}}$ is defined as the weighted L1 distance between the predicted depth sequence $\{\mathbf{d}_{1},\cdots,\mathbf{d}_{N}\}$ and ground truth depth $\mathbf{d}_{\text{GT}}$ with exponentially increasing weights following [lipson2021raft]:
[…]
where $\mu$ is set to 0.9. For the normal loss, given ground truth $\mathbf{N}_{\text{GT}}$, we combine the L1 distance and a perceptual loss [zhang2018unreasonable] to ensure both geometric accuracy and surface smoothness:
@@
We define the overall loss as a combination of $\mathcal{L}_{\text{albedo}}$, material smoothness loss $\mathcal{L}_{\text{smooth}}$, light transport loss $\mathcal{L}_{\text{LT}}$, and PBR loss $\mathcal{L}_{\text{PBR}}$ to facilitate appearance reconstruction and relighting via 2D-to-3D supervision:
@@
Note that after PBR and rasterization, the material maps: $\mathbf{A}\ \text{and}\ \boldsymbol{\Gamma}$; the visibility $V$ parameterized by SH map $\mathbf{Y}_{\text{vis}}$; the indirec light $\mathbf{L}_{\text{ind}}$ parameterized by SH map $\mathbf{Y}_{\text{ind}}$; the direct light $\mathbf{L}_{\text{d}}$, and the PBR color $\mathbf{C}_{\text{PBR}}$ are all rendered into the 2D image space as $\mathbf{I}_{\text{albedo}}$, $\mathbf{I}_{\text{rough}}$, $\mathbf{I}_{\text{ao}}$, $\mathbf{I}_{\text{indl}}$, $\mathbf{I}_{\text{dl}}$ and $\mathbf{I}_{\text{PBR}}$, respectively. Thus, they can be supervised via corresponding 2D ground truths. Specifically, the albedo loss is composed of an L1 loss and a perceptual loss to measure the difference between the predicted albedo and the ground truth $\mathbf{I}_{\text{albedo}}^{\prime}$:
@@
To promote spatially smooth material estimation, we adopt a bilateral smoothness loss following [gao2024relightable]:
$\begin{split}\mathcal{L}_{\text{smooth}}={}&\lambda_{3}\|\nabla\mathbf{I}_{\text{albedo}}\|\exp\left(-\|\nabla\mathbf{I}_{\text{albedo}}^{\prime}\|\right)\\ &+\lambda_{4}\|\nabla\mathbf{I}_{\text{rough}}\|\exp\left(-\|\nabla\mathbf{I}_{\text{albedo}}^{\prime}\|\right)\end{split}$ (23)
Note that ground truth roughness maps are not available; thus, the roughness map $\boldsymbol{\Gamma}$ is learned implicitly through the entire PBR framework. The light transport loss is combined with ambient occlusion loss $\mathcal{L}_{\text{ao}}$, direct light loss $\mathcal{L}_{\text{d}}$, and indirect light loss $\mathcal{L}_{\text{ind}}$, we use L1 distance to measure the predicted and ground truth ones:
@@
11 Limitations
While our method achieves high-quality relighting results on human subjects, it has two primary limitations. First, due to the intrinsic constraints of the adopted rendering equation, our approach cannot accurately model transparent or metallic materials, such as eyeglasses and metallic sequins, and struggles with extremely thin structures, including individual hair strands. Future work could explore the integration of advanced reflectance models to further improve the realism of relighting.
Second, our method requires at least two input images with precise camera poses for stereo-based geometry estimation and relighting. This dependency on multi-view calibration limits its applicability in scenarios with unknown or imprecise poses and reduces robustness in unconstrained environments. In future work, we aim to incorporate geometry priors from foundational models, such as VGGT [wang2025vggt], enabling novel view synthesis and high-quality relighting from arbitrary viewpoints without explicit calibration.
Figure 10: Dynamic relighting case from GPS-Gaussian [zheng2024gps] under diverse lighting conditions and novel viewpoints.
Figure 11: Dynamic relighting case from DNA-Rendering [cheng2023dna] under diverse lighting conditions and novel viewpoints.
Figure 12: Dynamic relighting case from DNA-Rendering [cheng2023dna] under diverse lighting conditions and novel viewpoints.
Figure 13: Dynamic relighting comparison on the GPS-Gaussian [zheng2024gps] dataset.
Figure 14: Dynamic relighting comparison on the DNA-Rendering [cheng2023dna] dataset.